# Ego-Exo4D: Understanding Skilled Human Activity from First- and Third-Person Perspectives

### v3

@@ 7.D Tools and Ecosystem
Jing Dong and Vijay Baiyya were responsible for obtaining camera poses, calibration, pointclouds and eye gaze using Aria MPS, created the 3D/4D visualizations for the paper and supplementary material, and acted as main contact points from the Aria team throughout the program; with Jing leading the algorithm development and verification, and Vijay leading the Aria MPS workflow and infrastructure development. Jakob Engel acted as technical and scientific advisor, and led the team that built the Aria Localization and Point Cloud algorithms. Kiran Somasundram helped design the capture setup and time-synchronization. Xiaqing Pan helped to align the Aria engineering team to support the EgoExo4D project. Mingfei Yan, Prince Gupta, and Sach Lakhavani acted as product managers of Aria and organizational leads for the successful use of Aria in the program. Kelly Forbes helped setting up agreements and working through the legal requirements of using Aria devices for recording the EgoExo4D dataset across the globe. Richard Newcombe initiated the Aria/Ego4D collaboration and acted as a scientific advisor throughout the program. Furthermore, we want to acknowledge the contribution of the entire Project Aria team as listed in [38], including Carl Ren and Sean Diener leading the Aria software and hardware engineering organization, and Renzo De Nardi as technical lead for the Aria device.
8   Camera Rig and Data Processing
The collection of ego-exo data at a global scale required us to develop a low-cost camera recording rig that was portable, auto-synchronized, and available internationally.
8.A Hardware
Our unified camera rig is as follows: 1 Aria, 4 GoPros [^4] This represents the common core of the collection rig used in all capture settings. In certain captures, additional exo or ego GoPros are also used, as noted in Appendix 10., 1 GoPro Remote, 4 Tripods, 4 SD Cards, 4 Tripod Mount Adapters, 4 Velcro’d Battery Packs, 4 USB-A to USB-C Cables, 1 Glasses Sports Strap, 1 Smartphone, 1 Laptop or Tablet for questionnaires. The total cost excluding the Aria/phone/laptop is under \$3,000, with the majority of that going to the GoPros.
8.B Time Sync
@@ 8.E Extensions
Contribution Statement
Procedural Physical
Cooking: Health: Music: Bouldering:
- Omelette - COVID test - Violin - V0 through V10
- Scrambled eggs - Cardiopulmonary - Piano
- Tomato and egg Resuscitation (CPR) - Guitar
- Sesame-ginger Asian salad
- Greek salad Bike repair: Basketball: Soccer:
- Sushi roll - Adjust rear
- Samosa derailleur (both limit Dance:
- Coffee latte screws & indexing) - Easy choreography
- Chai tea - Advanced choreography
- Milk
- Cookies
- Brownies
Table 1: Specific activities collected for the three Procedural and five Physical domains
Multimodal Egocentric Datasets
EGTEA-Gaze [81] 2018 V,A,G 32 1 7 106 15k - -
MECCANO [126] 2021 V,D,G 20 2 1 61 - - -
EK100 [29] 2022 V,A 37 45 N/A (97;300)* - - -
Ego4D [47] 2022 V,A,3D,S,G,I 931 74 N/A 110$\dagger$ - - -
Multiview Datasets
IXMAS [169] 2006 V 10 1 N/A 11 - - -
MEVA [26] 2021 V,T,GPS 100 28 N/A 37 - - -
Ego-Exo Datasets
CMU-MMAC [77] 2009 V,A,M,I 43 1 5 - - - -
Charades-Ego [144] 2018 1 71 N/A N/A 157 - - -
H2O [76] 2021 V,D 4 3 N/A 36 - - 0.5M
Assembly101 [139, 109] 2022 1 53 1 101 (24;90) - - 0.2M
EgoExo4D 2024 V,A,I,G,3D,6D,B,Ma 740 123 43 689 2.2M 9.6M 4.4M
[…]
9   Scenarios
Ego-Exo4D focuses on eight domains of human skill. See Table 1. These were selected to capture diversity, while also scoping our effort sufficiently to build a density of data within a core set of skills. Our scenarios were further targeted after a review of existing activity datasets, consideration of priority elements for transfer learning, and the identification of entry points to capture real-world scenarios available uniquely to our partners. To support data collection, we narrowed our scope to a small subset of tasks within each domain.
As discussed in the paper, we divide these into the categories of procedural activities—requiring a sequence of steps to achieve a goal state—and physical activities—requiring some degree of refined physical control to perform appropriately, but not necessarily following a step-by-step order. Note that in general an activity can be both procedural and physical; they are not mutually exclusive properties. Each consortium member selected multiple domains of focus from the eight targeted domains (cooking, health, etc.) based on their own preferences and local opportunities to capture data of these scenes at scale.
In working groups, our consortium developed data collection guidelines for each of our eight domains. These guidelines described the recommended camera positioning, instructions for participating camera wearers, along with important context-specific considerations. For example, given privacy concerns, our health guidelines required data collection participants to discard COVID tests before results were visible. These guidelines provided general parameters from which to collect data; however, they were not rigid steps.
To support diversity and implementation at a global scale, we allowed for site-specific adjustments to be made. For this reason, our dataset contains important nuances that should be noted, including, for example, differing standards for the implementation of CPR, cultural differences in the ingredients used for different targeted dishes, and location-specific bouldering routes.
Additionally, while data collection partners were encouraged to select scenarios of their preference, we did set broad targets for each site to collect at least three domains among the eight options. Since every domain is covered by more than one partner, the dataset exhibits visual variety from the different physical locations. Even within the captures done by a single partner, there are often multiple different sites used for filming (e.g., a couple different bike shops in the same city).
These domain selections were additional to our one cross-cutting scenario of cooking, which was collected by all partners. Cooking was a priority domain because it resonates around the world as a human need and interest. We are pleased to report that the cooking scenario of Ego-Exo4D contains more than 650 takes of cooking performed by more than 170 chefs in 60 different environments around the world, forming nearly 100 hours of ego video alone.
Figure 13 and 14 shows example frames illustrating the variety of the sites and tasks.
Figure 13: Ego-Exo4D captures skilled activity from 8 domains, in a wide variety of 123 scenes in 13 different cities in Japan, Colombia, Canada, India, Singapore, and seven US states. Every odd column shows an ego view, and the adjacent even column shows one of its paired exo views.
Figure 14: Ego-Exo4D captures skilled activity from 8 domains, in a wide variety of 123 scenes in 13 different cities in Japan, Colombia, Canada, India, Singapore, and seven US states. Every odd column shows an ego view, and the adjacent even column shows one of its paired exo views.
Contribution Statement
Bike Repair  Yale Song led the scenario development and collection guidelines for Bike Repair. Other contributors in the working group were David Crandall and Jim Rehg.
Culinary  Manolis Savva led development of the scenario and collection guidelines. Other contributors in the working group: Andrew Westbury, Kumar Ashutosh, Deepti Ghadiyaram, Gene Byrne, Kristen Grauman, Santhosh Kumar Ramakrishnan, Shun Iwase, Yan Xu
Health  Mike Zheng Shou was the lead author of the health scenarios guidelines.
Bouldering and Dance  Pablo Arbeláez and Maria Escobar were the lead authors of the Bouldering and Dance guidelines.
Music Jianbo Shi and the team from the University of Pennsylvania were the lead authors of the Music guidelines.
Basketball  Gedas Bertasius was the lead author of the Basketball guidelines.
Soccer Rawal Khirodkar was the lead author of the Soccer guidelines, camera placement, size of the capture area, and drills that were captured.
Refer to caption Refer to caption
10   Data Collection
@@
10.A   Carnegie Mellon University
@@
10.B   FAIR, Meta
We collected 119 total takes of skills demonstrations in New York and three different locations in California. We focused on cooking and bike repair, taking advantage of the skilled workforce of chefs and bike technicians that serve our campuses and employees in these locations. We used the unified camera rig of 1 Aria and 4 GoPros without any additional sensors.
@@ 10.B   FAIR, Meta
Cooking Our chefs recorded five different recipes as part of 17 unique takes, including salads, egg dishes, and Asian Garlic Noodles. Locations featured in the dataset are three different professional kitchens used to prepare and serve hundreds of employees each day.
@@
In total, we were able to mobilize five chefs and four bike mechanics. Due to workplace considerations, we did not collect demographic or age information about our chef and mechanic partners. Participating chefs and bike technicians are highly skilled with all research subjects reporting that they do the activity shown in the dataset daily or weekly. Similarly, eight research subjects have more than 10 years professional experience.
10.C   Georgia Tech
@@ 10.E   Indiana University
10.F   National University Singapore
@@
10.G   Simon Fraser University
@@ 10.H   Universidad de los Andes
10.I   University of Minnesota
Collection at the University of Minnesota, Twin Cities focused on two main scenarios: Bouldering and Cooking. A total of 249 takes with 53 unique participants were collected. We collected all data using the unified camera rig with no additions.
@@ 10.I   University of Minnesota
10.J   University of North Carolina
@@
Basketball All participants performed three basketball drills: Mikan Layup, Reverse Layup, and Mid-range Jumpshooting, for 388 takes. We recruited 11 expert players from the university team with 10+ years of experience. To improve the participant skill diversity in the dataset, we also recruited novice players with less than 1 year of playing experience. The location of data collection was a university basketball gym.
Music For the music scenario, we asked all 9 participants to play 5 minutes of scales and arpeggios and 10 minutes of free play for 27 takes. All of our participants were recruited from the university music club and considered themselves as experts at playing their respective instruments. The instruments featured in our collected dataset were piano, trombone, trumpet, and saxophone. The Ego-Exo4D music guidelines called for just piano, violin, and guitar, but we found it necessary to expand this list in order to gather data for this domain. All data was recorded in a university music room.
Soccer For soccer, we focused on three drills: dribbling, juggling, and penalty shots for 133 takes across 12 unique participants. 7 of these participants were experts with 10+ years of experience, whereas the remaining 5 participants were casual soccer players. All videos were collected at a university soccer field.
@@ 10.K   University of Pennsylvania
Cooking, categorized as "procedural", consisted of preparing four dishes: an egg dish, a salad, a noodle dish, and a dessert. The group of participants consisted primarily of Penn students with experience ranging from amateurs to hobbyists. Professionals were unavailable due to scheduling conflicts. We totaled 81 takes over 20 participants. The entire filming process was undertaken within a three-week span, primarily at the apartment of one of the team’s participants. This location expedited our data collection for this task by providing a stove and fridge for regular use.
@@
Dance captures, classified as a "physical", consist of four takes of dancers performing dance routines to a song. The dance types recorded included Lindy-Hop Jazz, Bollywood, Latin, and Chinese Folk Dance; across these genres, we totaled 165 takes over 38 participants. The Lindy-Hop Jazz dancers came from the Jazz Swing Attacks, a dance club in Philadelphia. Contact was established via Instagram, and data, collected weekly over a month. This group contained a balanced mix of experienced instructors and beginner dancers. The Bollywood dancers, the Drexel’s Philly Maza, were recorded in the Drexel Engineering Building. They compete nationally but routinely train beginner recruits. The Chinese Folk Dancers were members of the local Great Wall Chinese School’s dance club and independent student volunteers with prior competitive dance training. These were captured in the SIG Lab for collections.
@@ 10.L   University of Tokyo
Cooking We recruited 12 Japanese participants living in the Tokyo area through a temporary employment agency. The gender and age of the participants were balanced to collect diverse behavior patterns. The participants cook three days or more each week in their daily lives. Each participant prepared three dishes: an omelet, a white radish & lettuce & tomato & cucumber salad, and a sushi roll. We recorded both versions with and without narrations for each dish and participant. A one-page summary of each recipe was provided before data collection and was shown during video recording so that the participants could prepare the dishes smoothly, and the procedure of each recipe should be consistent between the participants.
@@
Health We recruited 17 Japanese participants living around Tokyo, Japan, through a temporary employment agency. The gender and age of the participants were balanced to collect diverse behavior patterns. We recorded videos of the 17 participants performing two tasks: COVID-19 rapid antigen testing using three test kits and performing CPR on a mannequin. We conducted all recordings in the same meeting room on campus over two days. For the COVID test, an instruction manual of each test kit was provided to the participants before and during the recording. Also, we did not show the participants or record any COVID test results for privacy protection. For CPR, the participants took an introductory lifesaving course provided by the Tokyo Fire Department before recording. Besides, a one-page summary of the CPR procedure was provided before the data collection and shown during the video recording. This is so that the participants can perform CPR smoothly and the procedure is consistent among the participants. We collected 73 takes from 17 participants. All of the CPR procedures (17 takes) were recorded without narration. For the COVID test, we recorded the videos of the 17 participants using the three test kits (51 takes) without narrations. The video length of each participant’s performing CPR is 9 min 48 sec on average, including the camera synchronization procedure. Similarly, the video length of each participant’s using the three COVID test kits is 29 min 22 sec on average. Additionally, we recorded extra takes of 5 participants out of the 17 using a test kit with narrations (5 takes in total).
@@
Contributions Statement
Carnegie Mellon University Rawal Khirodkar developed the automatic 3D body keypoints extraction pipeline and collected a subset of the soccer, bike mechanic, and cooking sequences for the CMU portion of the dataset. Sean Crane was in charge of the data collection, IRB documents, capturing data, working with participants and processing the data for CMU. Abrham Gebreselasie ran the actionformer-based baseline for demonstration proficiency benchmark. Eugene Byrne served as the engineering lead for the initial design and implementation of the dataset, camera rig, processing pipeline and keystep annotations while at Meta. Subsequently at CMU, he assisted in the recognition benchmarks, implemented Ego-Exo transfer  [82] (1 of the 3 baseline methods for keystep recognition) and the initial implementation of keystep action detection  [189], and assisted in annotation/data quality generally.
Indiana University David Crandall oversaw the overall effort at Indiana University, including protocol design and data collection. At Indiana University, Weslie Khoo led the IRB protocol design and compliance, arranged logistics such as ordering equipment, and designed and oversaw the cooking scenario data collection. Yuchen Wang and Ziwei Zhao co-led the participant recruitment and data collection for all three scenarios. Ziwei Zhao led data preparation and transfer. We also acknowledge Manasi Swaminathan who assisted with data collection and video synchronization.
11   Participants
Next we overview the background of the participants (camera wearers) in the Ego-Exo4D dataset.
11.A   Demographics
We provide self-declared information on ethnic groups by the participants. Sharing this information was optional for all research subjects. Ethnicity is reported based on location specific categories as defined by the relevant partner lab. No such information was gathered from research subjects participating in our collections in California, New York, and Pittsburgh, Pennsylvania. Aggregated gender and age information of all participants in Ego-Exo4D is provided in Figure 20.
Figure 20: Participants’ self-reported demographic information
@@ 11.A   Demographics
11.B   Participant Surveys
To ensure consistent, high quality annotations (which will be discussed in Appendix 13), we identified three aspects from data collection where additional information from the participants’ recording is crucial. These are: the skill of the participant, the objects they are using, and the actions they are completing. The three aspects represent information that either cannot be captured by other annotators after the fact or could result in noise within the annotation process. Accordingly, we designed participant surveys to cover key information regarding the participants’ skill. To capture the actions and objects with which they interact, we ask participants to perform a round of first-person narrations called “narrate-and-act" (cf. Appendix 12.B).
@@ 11.B   Participant Surveys
Pre-task Questionnaire
Within this questionnaire, the aim is to capture the participant’s perceived skill level regarding the task that they are about to perform. To maximize consistency across participants, we forgo asking participants to self-rate themselves as experts, novice, etc.—instead, we ask participants more quantifiable questions such as length of time the participant has been doing this task and frequency. [^5] We also obtain proficiency ratings for the participants via our expert commentators (cf. Appendix 12.A). Additional questions regarding whether the participant has taught the task to others and whether they have recorded/watched videos about the task also give signals regarding the proficiency of the user. Note that the information gathered is scenario- and take-specific, since the number of hours to be considered an expert or amateur can be wildly different (e.g. cooking vs. bouldering).
Post-task Questionnaire
The post-task questionnaire focuses on capturing details of the take and how it went according to the participant. Questions are asked about mistakes, objects, and time taken compared to predicted time, which is beneficial for downstream annotation.
The survey data will be shared with the dataset.
@@
Figure 21: Examples of the three different annotation styles: narrate and act, atomic action descriptions, and expert commentaries from four of the scenarios (bike repair, health, dance, and basketball). We also include word clouds which highlight the differences in vocabularies per scenario. In narrate and act text, we see how the participants briefly describes what they are doing and why, whereas the atomic action descriptions provide strictly a statement about the visible actions. The expert commentary offers an expert’s critique of what is shown, commenting on strengths and weaknesses and explaining how the participant’s actions affect their performance.
12   Language Descriptions
As introduced in the main paper, Ego-Exo4D provides three forms of parallel text corpora for the video: expert commentary, narrate and act, and atomic action descriptions. Figure 21 and Table 4 show examples from different scenarios highlighting their distinctions in style and point of view. In the following we elaborate on their protocols, show examples, and analyze their differences.
Ego-Exo4D’s language annotations are not targeted at any particular benchmark task or taxonomy. They aim to be a general resource that will inspire new language-vision possibilities, such as learning how to generate coaching or teaching advice given a demonstration video, learning how to spot skill and lack of skill, or learning general multimodal embeddings in the procedural and skilled activity domain. We also anticipate the temporally grounded descriptions to be valuable for pre-training foundation models [85, 123] or automated video captioning [111, 59, 199]. Furthermore, the time-anchored aspect of the three language annotations provides the opportunity to retrieve time points in specific Ego-Exo4D videos that correspond to queried moments, actions, or phrases. Finally, they are valuable to mine the dataset for the distribution of objects and activities present, e.g., for taxonomy formation.
12.A   Expert Commentary
Ego-Exo4D offers a new video-language resource specifically designed to support the study of human skill. We mobilize 52 expert commentators, each with rich qualifications and skills in one of our focus domains, to provide time-anchored critiques and explanations of all the skill demonstrations in the dataset. Specifically, our expert commentators use spoken word to describe what is most effective or ineffective about the camera wearer’s actions, review the quality of the execution, and identify mistakes. Experts also provide an overall proficiency rating on each skills demonstration, assessing how well the task was performed with a short written justification. The following section describes the qualifications and background of our experts; commentary instructions and tooling; and statistics on collected commentaries to date.
Experts and Expert Qualifications
Over a period of three months, we recruited experts from across each of Ego-Exo4D’s eight core domains, working with a selection panel of ten individuals to review qualifications and engage in more than 68 interviews. Our selection criteria focused on technical skills, communication, and task completion along with performance reviews during a live video commentating exercise. Among the numerous candidates we interviewed, we were pleased to select 52 to join the effort.
On average, 90% of recruited experts possess more than 10 years of professional experience and all have served during this time in the capacity of a coach, instructor or mentor. All experts further have either an advanced degree in their domain of focus or an industry certification. Certification authorities include the US Soccer Federation, the American Culinary Federation, USA Climbing, the American Red Cross, Trek Bikes, and New York State’s Initial Certification in Teaching Dance, among others.
Multiple individuals were recruited across each domain, with the goal of generating language and expertise diversity. Table 5 shows the number of expert contributors per domain. Due to employment considerations, all experts are residents of the United States.
Figure 22: Two examples of expert commentary and proficiency scores, along with spatial drawings (red) done by the expert to augment the spoken comments.
Commentary Instructions and Tooling
Experts are provided with two time-synchronized videos of each Ego-Exo4D skills demonstration—one showing the egocentric view and another providing a single exocentric perspective specifically selected by annotators as the view that provides the best visibility on the scene. Experts are then asked to watch the video in full without commenting to gain an understanding of the skills demonstration and plan out important points to note in their commentary.
After this initial viewing, experts are asked to provide a numeric proficiency score on the scale of 1 (least skilled) to 10 (most skilled), rating how well the camera wearer performed the task. A short written description of why the chosen score was selected is also provided at this stage. In many cases, experts coordinated within their domain group to calibrate this scoring. While each proficiency rating is at the discretion of individual experts, domain-level calibration provided a general framework for assessing skills demonstrations.
With proficiency scores noted, experts play the video from the start and pause at critical moments of the demonstration to record verbal descriptions of what is observed to be most effective or ineffective about the camera wearer’s actions, offer performance tips, identify mistakes or suboptimal execution, and why they matter. In all cases, commentary is time-anchored and retrospective, focusing on insights and perspectives relating to actions visible to that point in the video. We choose to collect commentary as verbal recordings in order to maintain the naturalness of the performance descriptions and do so quickly. In all cases, experts had the option to use a “telestrator” tool to enhance their commentary with freehand sketches to spatially localize information or otherwise help explain a point. See Figure 22. We transcribe the commentaries automatically with OpenAI’s Whisper for automatic speech recognition [125]. Table 4 (right column) and Figure 21 shows example commentaries. In total, our experts have dedicated more than 6,000 hours of effort to date to provide this commentary.

### v4

@@ 6 Conclusions
Acknowledgements
We gratefully acknowledge the following colleagues for valuable discussions and support of our project: Vittorio Caggiano, Sarah Carroll, Ilé Danza, Ahmad Darkhalil, Zona de Bloque, Alex Dinh, Rene Martinez Doehner, Ivan Cruz, Matt Feiszli, Vance Feutz, Kelly Forbes, Rohit Girdhar, Pierre Gleize, Andrés Hernández, Shun Iwase, Hanxiao Jiang, Armin Kavian, Bolin Lai, Vivian Lee, Brighid Meredith, Ashley Massie, Natalia Neverova, Manohar Paluri, Joelle Pineau, Artsiom Sanakoyeu, Paresh Shenoy, Jiaray Shi, Jiasheng Shi, Gaurav Shrivastava, Mitesh Singh, Manasi Swaminathan, Arjang Talattof, Ali Thabet, Laurens van der Maaten, Andrea Vedaldi, and Tobby Zhu.
We also sincerely thank the 52 experts who contributed to the expert commentary for their expertise and support; they are listed individually in Appendix D. Thank you to the Common Visual Data Foundation (CVDF) for hosting the Ego-Exo4D dataset. Finally, thank you to the 740 participants who contributed to this dataset and shared their skills in video.
UT Austin is supported in part by the IFML NSF AI Institute. University of Catania is supported in part by the project Future Artificial Intelligence Research (FAIR) – PNRR MUR Cod. PE0000013 - CUP: E63C22001940006. Luigi Seminara is supported by PNRR PhD scholarship “Digital Innovation: Models, Systems and Applications” DM 118/2023. Simon Fraser University is supported in part by the Canada Research Chairs Program (CRC-2019-00298) and NSERC Discovery (2019-06489) Georgia Tech is supported in part by NSF (#2144194) and NSF-GRFP Indiana University is supported in part by NSF DRL-2112635 (AI Institute for Engaged Learning). Univ. of Bristol is supported in part by EPSRC UMPIRE (EP/T004991/1) and EPSRC PG Visual AI (EP/T028572/1). Z. Zhu is supported by UoB-CSC Scholarship. University of Tokyo is supported in part by JST ASPIRE Grant Number JPMJAP2303 and JSPS KAKENHI Grant Numbers JP24K02956.
Appendix A Camera setup and recording details
A.1 Time sync
@@ A.3 Recording procedure
A.4 Aria post-processing
First, the Aria Machine Perception Services (MPS) pipeline is invoked for each full Aria recording—these typically are about 20 minutes to 1 hour long and can include several takes, the hand-over in-between takes, as well as some other set-up steps. This is followed by localizing all GoPro videos of that scene as described above, and finally followed by time-synchronization across Aria and the GoPro cameras, as well as take-separation, as described below.
Appendix B Data collection
@@
B.1 Carnegie Mellon University
@@
B.2 FAIR, Meta
We collected 119 total takes of skills demonstrations in New York and three different locations in California. We focused on cooking and bike repair, bringing in a skilled workforce of chefs and bike technicians that serve major kitchens and repair shops in the area. We used the unified camera rig of 1 Aria and 4 GoPros without any additional sensors.
@@ B.2 FAIR, Meta
Cooking Our chefs recorded five different recipes as part of 17 unique takes, including salads, egg dishes, and Asian garlic noodles. Locations featured in the dataset are three different professional kitchens used to prepare and serve hundreds of people each day.
@@
In total, we were able to mobilize five chefs and four bike mechanics. Participating chefs and bike technicians are highly skilled, with all research subjects reporting that they do the activity shown in the dataset daily or weekly. Similarly, eight research subjects have more than 10 years professional experience.
B.3 Georgia Tech
@@ B.5 Indiana University
B.6 National University Singapore
@@
B.7 Simon Fraser University
@@ B.8 Universidad de los Andes
B.9 University of Minnesota
Collection at the University of Minnesota focused on two main scenarios: Bouldering and Cooking. A total of 249 takes with 53 unique participants were collected. We collected all data using the unified camera rig with no additions.
@@ B.9 University of Minnesota
B.10 University of North Carolina
@@
Basketball
All participants performed three basketball drills: Mikan Layup, Reverse Layup, and Mid-range Jumpshooting, for 388 takes. We recruited 11 expert players from the university team with 10+ years of experience. To improve the participant skill diversity in the dataset, we also recruited novice players with less than 1 year of playing experience. The location of data collection was a university basketball gym.
Music
For the music scenario, we asked all 9 participants to play 5 minutes of scales and arpeggios and 10 minutes of free play for 27 takes. All of our participants were recruited from the university music club and considered themselves as experts at playing their respective instruments. The instruments featured in our collected dataset were piano, trombone, trumpet, and saxophone. The Ego-Exo4D music guidelines called for just piano, violin, and guitar, but we found it necessary to expand this list in order to gather data for this domain. All data was recorded in a university music room.
Soccer
For soccer, we focused on three drills: dribbling, juggling, and penalty shots for 133 takes across 12 unique participants. 7 of these participants were experts with 10+ years of experience, whereas the remaining 5 participants were casual soccer players. All videos were collected at a university soccer field.
@@ B.11 University of Pennsylvania
Cooking, categorized as ”procedural”, consisted of preparing four dishes: an egg dish, a salad, a noodle dish, and a dessert. The group of participants consisted primarily of Penn students with experience ranging from amateurs to hobbyists. Professionals were unavailable due to scheduling conflicts. We totaled 81 takes over 20 participants. The entire filming process was undertaken within a three-week span, primarily at the apartment of one of the team’s participants. This location expedited our data collection for this task by providing a stove and fridge for regular use.
@@
Dance captures, classified as a ”physical”, consist of four takes of dancers performing dance routines to a song. The dance types recorded included Lindy-Hop Jazz, Bollywood, Latin, and Chinese Folk Dance; across these genres, we totaled 165 takes over 38 participants. The Lindy-Hop Jazz dancers came from the Jazz Swing Attacks, a dance club in Philadelphia. Contact was established via Instagram, and data, collected weekly over a month. This group contained a balanced mix of experienced instructors and beginner dancers. The Bollywood dancers, the Drexel’s Philly Maza, were recorded in the Drexel Engineering Building. They compete nationally but routinely train beginner recruits. The Chinese Folk Dancers were members of the local Great Wall Chinese School’s dance club and independent student volunteers with prior competitive dance training. These were captured in the SIG Lab for collections.
@@ B.12 University of Tokyo
Cooking
We recruited 12 Japanese participants living in the Tokyo area through a temporary employment agency. The gender and age of the participants were balanced to collect diverse behavior patterns. The participants cook three days or more each week in their daily lives. Each participant prepared three dishes: an omelet, a white radish & lettuce & tomato & cucumber salad, and a sushi roll. We recorded both versions with and without narrations for each dish and participant. A one-page summary of each recipe was provided before data collection and was shown during video recording so that the participants could prepare the dishes smoothly, and the procedure of each recipe should be consistent between the participants.
@@
Health
We recruited 17 Japanese participants living around Tokyo, Japan, through a temporary employment agency. The gender and age of the participants were balanced to collect diverse behavior patterns. We recorded videos of the 17 participants performing two tasks: COVID-19 rapid antigen testing using three test kits and performing CPR on a mannequin. We conducted all recordings in the same meeting room on campus over two days. For the COVID test, an instruction manual of each test kit was provided to the participants before and during the recording. Also, we did not show the participants or record any COVID test results for privacy protection. For CPR, the participants took an introductory lifesaving course provided by the Tokyo Fire Department before recording. Besides, a one-page summary of the CPR procedure was provided before the data collection and shown during the video recording. This is so that the participants can perform CPR smoothly and the procedure is consistent among the participants. We collected 73 takes from 17 participants. All of the CPR procedures (17 takes) were recorded without narration. For the COVID test, we recorded the videos of the 17 participants using the three test kits (51 takes) without narrations. The video length of each participant’s performing CPR is 9 min 48 sec on average, including the camera synchronization procedure. Similarly, the video length of each participant’s using the three COVID test kits is 29 min 22 sec on average. Additionally, we recorded extra takes of 5 participants out of the 17 using a test kit with narrations (5 takes in total).
@@
Appendix C   Participants
We provide self-declared information on ethnic groups by the participants. Sharing this information was optional for all research subjects. Ethnicity is reported based on location specific categories as defined by the relevant partner lab. No such information was gathered from research subjects participating in our collections in California, New York, and Pittsburgh, Pennsylvania.
@@ Appendix C   Participants
Appendix D Language Descriptions